OpenStudy (anonymous):

Given the equation y - 3 = 1/2(x + 6) in point-slope form, identify the equation of the same line in standard form.

sammixboo (sammixboo):

What???

OpenStudy (anonymous):

Nevermind

sammixboo (sammixboo):

?????

OpenStudy (anonymous):

I asked a question earlier, and you posted to it: "Welcome to open study, if you have any questions, ask me" so, I asked.

OpenStudy (anonymous):

@satellite73 can you help me please?

sammixboo (sammixboo):

What did you ask??? If that was about an hour ago that was on my backup account...

OpenStudy (anonymous):

are you trying to write \[y-3=\frac{1}{2}(x+6)\] in standard form?

OpenStudy (anonymous):

I was about an hour ago, yes. @satellite73 Yes.

OpenStudy (anonymous):

first step would be to get rid of that annoying \(\frac{1}{2}\) by multiplying everything by \(2\) and get \[y-6=x+6\]

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

oops typo there \[2y-6=x+6\] is more like it

OpenStudy (anonymous):

Haha, I knew what you meant.

OpenStudy (anonymous):

man i am full of typos lets try that again starting with \[2y-6=x+6\] SUBTRACT \(2y\) and get \[-6=x-2y+6\] and then subtract \(6\) to get \[-12=x-2y\]

OpenStudy (anonymous):

XD Thank you so much!

OpenStudy (anonymous):

yw